OpenStudy (anonymous):

x^4 - 5x^2 - 14 = 0 substitute u for x^2 I dont understand how to solve this one

OpenStudy (anonymous):

x^4 is really (x^2)^2 so you would end up with u^2-5u-14=0

hero (hero):

\[(x^2)^2 - 5x^2 - 14 = 0\]\[x^2 = u\]\[u^2 - 5u - 14 = 0\]

OpenStudy (anonymous):

awesome thank you. I was confused

OpenStudy (campbell_st):

its simply becomes a quadratic \[u^2 - 5u - 14 = 0\] solve by factoring then let x^2 = the solutions you get for u

OpenStudy (anonymous):

Great thank you very much ^_^

OpenStudy (anonymous):

If u = x^2, then u^2 = x^4 So your equation is just u^2 -5u - 14 = 0 Factor this, (u-7)(u+2) = 0 so, u = 7 and u = -2 Now, since u = x^2, so x^2 = 7 , then x = +/- sqrt(7) If u = - 2, then x^2 = -2, x = +/- sqrt (-2) = +/- i sqrt(2)